OpenStudy (anonymous):

help please? Find the value of the annuity. a[1]=$3000, I=0.06, n=20

OpenStudy (mertsj):

What is the formula?

OpenStudy (anonymous):

S[n]=a[1] AND inside brackets is (1+I)^n-1 OVER I

OpenStudy (anonymous):

the best way I can write the formula

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

\[s _{n}= a _{1}(1+I)^{\frac{n-1}{I}}\]

OpenStudy (mertsj):

So is that the correct formula?

OpenStudy (anonymous):

its that formula, but the line is inside the parenthesis with the i on the bottom

OpenStudy (mertsj):

\[S _{n}=3000(1+.06)^{\frac{20-1}{.06}}\]

OpenStudy (mertsj):

I guess I don't know what you mean when you say the line is inside the parentheses.

OpenStudy (mertsj):